OpenStudy (abbycross167):

Can someone please help me answer this question? Please? I would really appreciate it... Question~ Solve the systems using the elimination method. y = 2x - 1 y = -x + 7

OpenStudy (abbycross167):

My Answer~ Y=x-1A -2x+y=-1 X+y=7 = -2x+y=-1 Multiply by 1 X+y=7 multiply by 2 = 2x-1=-x+7 3x-1=7 3x=8 X=8/3 Y=13/3 Final answer~ (8/3, 13,3)

OpenStudy (abbycross167):

What my teacher said was wrong~ -2x+y=-1 x+y=7 good! Yes, multiply by 2, but you didn't do it!?!

OpenStudy (abbycross167):

Can someone please show me how to fix my answer?

OpenStudy (photon336):

sure

OpenStudy (photon336):

\[y = 2x-1 \] \[y_{2} = -x+7\] we could easily eliminate y by multiplying the top equation by -1 and keeping the bottom the same. so we get this. then we combine them and find that -y+y = 0 and we're left with x. -y = -2x+1 y = -x+7 0 = -3x+8 -8 = -3x x = (8/3) now we have x = (8/3) the next step is easier. to find y we can plug this in -(8/3)+7 = y y = (13/3)

OpenStudy (photon336):

@abbycross167 yep that's correct.

OpenStudy (abbycross167):

@Photon336 why is y = -x + 7(from my question) written as y2= -x + 7 in your answer?

OpenStudy (abbycross167):

@Photon336

OpenStudy (photon336):

@abbycross167 just ignore that lol that was a typo

OpenStudy (abbycross167):

ooh ok haha that you so much for helping me!!